# The Radio Counterpart to the Fast X-ray Transient EP240414a

### v1

@@ Abstract
Despite being operational for only a short time, the Einstein Probe mission has already significantly advanced the study of rapid variability in the soft X-ray sky. We report the discovery of luminous and variable radio emission from the Einstein Probe fast X-ray transient EP240414a, the second such source with a radio counterpart. The radio emission at $3\,\rm{GHz}$ peaks at $\sim 30$ days post explosion and with a spectral luminosity $\sim 2\times 10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$, similar to what is seen from long gamma-ray bursts, and distinct from other extra-galactic transients including supernovae and tidal disruption events, although we cannot completely rule out emission from engine driven stellar explosions e.g. the fast blue optical transients. An equipartition analysis of our radio data reveals that an outflow with at least a moderate bulk Lorentz factor ($\Gamma\gtrsim 1.6$) with a minimum energy of $\sim 10^{48}\,\rm{erg}$ is required to explain our observations. The apparent lack of reported gamma-ray counterpart to EP240414a could suggest that an off-axis or choked jet could be responsible for the radio emission, although a low luminosity gamma-ray burst may have gone undetected. Our observations are consistent with the hypothesis that a significant fraction of extragalactic fast X-ray transients are associated with the deaths of massive stars.
@@ 1 Introduction
The observational paradigm for FXTs has recently undergone a drastic shift with the launch of the Einstein Probe X-ray telescope (EP; Yuan et al., 2022). The two instruments on board the EP, the Wide Field X-ray Telescope (WXT) and the Follow-up X-ray Telescope (somewhat unfortunately abbreviated as FXT), allow for both wide sky area monitoring as well as follow-up and localisation regions that range between arcminutes and tens of arcseconds, with FXT candidates reported via public alert streams. These capabilities have allowed for rapid follow-up at optical, radio, and X-ray wavelengths and the detection of multi-wavelength counterparts to EPW20240219aa, EP240305a, EP240309a, and EP240315a. The sources EP240305a and EP240309a are likely Galactic, and have been associated with variability from a Gaia star (Liu et al., 2024a) and a cataclysmic variable (Buckley et al., 2024), respectively. EPW20240219aa was associated with a sub-threshold event by the Fermi Gamma-ray Burst Monitor and is therefore thought to have been caused by a gamma-ray burst (Fletcher et al., 2024). EP240315a was associated with a high redshift ($z\sim 4.9$) galaxy and was followed up across the electromagnetic spectrum, being the first EP FXT with a radio and optical counterpart (Carotenuto et al., 2024; Gillanders et al., 2024). Analysis of the radio, optical, and X-ray data lead to the conclusion that the event harboured a relativistic jet, and was likely from a long gamma-ray burst (Gillanders et al., 2024; Ricci et al., 2024; Levan et al., 2024b; Liu et al., 2024b). With these events, and the increasing number of new events being discovered discovered by the EP, it is clear that significant progress can now be made in understanding the progenitors to FXTs and their multiwavelength properties.
@@
In this paper we present a radio observing campaign on EP240414a, the second EP extragalactic FXT with a multi-wavelength counterpart. EP240414a was discovered (Lian et al., 2024) by the WXT at UTC 09:50:12 2024-04-14 (MJD 60414.4099; which we define to be $T_{0}$) with a peak flux of $\sim 3\times 10^{-9}\,\rm{erg}\,\rm{s}^{-1}\,\rm{cm}^{-2}$ in the $0.5$-$4\,\rm{keV}$ energy band and was seen to fade by 4 dex over the following 2 hours with the EP Follow-up X-ray Telescope (Guan et al., 2024). An optical counterpart (AT2024gsa) was discovered within three hours of the FXT detection (Aryan et al., 2024) and a redshift of the nearest galaxy was reported to be $z=0.41$ (Jonker et al., 2024). The rapid and unusual optical lightcurve of the source is reported in Sirvastav et al. (2024) with the redshift confirmed to be $z=0.4018\pm 0.0010$ (see also van Dalen et al. 2024). This implies a $\sim 27\,\rm{kpc}$ projected offset from the host and a peak X-ray luminosity of $\sim 10^{48}\,\rm{erg}\,\rm{s}^{-1}$, similar to those seen from GRB afterglows at early times (see e.g. Gehrels et al., 2009). We discovered the radio counterpart to EP240414a with the MeerKAT radio interferometer, six days after it was reported as an FXT by the EP, as a $\sim 200\,\mu\rm{Jy}$ radio source coincident with the EP error circles (Guan et al., 2024; Lian et al., 2024) and coincident with AT2024gsa (Bright et al., 2024; Aryan et al., 2024, Srivastav et al. 2024; van Dalen et al. 2024).
@@ 2.1 MeerKAT
Figure 1: Optical and radio detections of EP240414a/AT2024gsa. (top) A Pan-STARRS i-band image of the field of EP240414a. The source position is marked by a pair of white lines. The diffuse source to the left is the putative host galaxy SDSS J124601.99--094309.3 (Aryan et al., 2024; Jonker et al., 2024) at a redshift of $z=0.4018\pm 0.0010$ (Srivastav et al. 2024, see also van Dalen et al. 2024). (bottom) A subsection of our MeerKAT discovery image of EP240414a with the source position marked by a pair of black lines. The source has a flux density of $227\pm 13\,\mu\rm{Jy}$. We detect clear emission from the possible host galaxy of EP240414a. The MeerKAT restoring beam is shown in the bottom left and is $3.2^{\prime\prime}\times 3.2^{\prime\prime}$ at a position angle of 0 degrees. A scale bar in the bottom right shows $10\,\rm{kpc}$ at $z=0.4$.
@@ 2.2 Australia Telescope Compact Array
We obtained radio observations of EP240414a with the Australia Telescope Compact Array (ATCA) under program CX576 (PI Carotenuto). We first observed EP240414a on 2024 July 2 between 06:10 UT and 10:11 UT. ATCA was in its extended 6D configuration. A second observation, under the same program, was performed on August 23rd 2024 between 04:40 and 08:30 UT, with the telescope in the more compact 1.5A configuration. For both epochs, data were recorded simultaneously at central frequencies of 5.5 GHz and 9.0 GHz, with 2 GHz of bandwidth at each frequency. We used PKS B1934--638 for bandpass and flux density calibration, and J1239–1023 for the complex gain calibration. Data were flagged, calibrated, and imaged using standard procedures within CASA. When imaging, we used a Briggs robust parameter of 0 to balance sensitivity and resolution. Our ATCA observations are summarised in Table 1.
@@ 2.3 Allen Telescope Array
As part of a larger survey of EP transients we began observing the field of EP240414a on 2024 May 5 with the Allen Telescope Array (ATA; see Farah et al. in prep.; Pollak et al. in prep.). While multiple observing bands were used (see e.g. Bright et al., 2022, for details) we report only our most constraining limit at a central frequency of $8\,\rm{GHz}$. Observations were reduced with a custom pipeline utilising CASA for calibration and wsclean for imaging. 3C286 was used to set the absolute flux scale and bandpass response of the instrument, while 1246--075 was used to correct for the time-dependent complex gains. Our ATA observations are summarised in Table 1.
@@ 3 Results
Our radio observations of EP240414a are fairly sparse, with three observations at $3\,\rm{GHz}$ with MeerKAT, and two ATCA observations both at $5.5$ and $9\,\rm{GHz}$ (see Table 1 and Figure 2). Our $3\,\rm{GHz}$ light curve shows a clear peak at around $\sim 30$ days post explosion in the observer reference frame (or $\sim 20$ days in the source rest frame) at $434\pm 21\,\mu\rm{Jy}$. While our sampling rate is not high enough to confidently confirm this as the exact peak time at $3\,\rm{GHz}$, we can make the approximation that a spectral peak moved through $3\,\rm{GHz}$ at $30$ days post explosion and therefore take $T_{\rm{pk},3\,\rm{GHz}}=30$ days and $F_{\rm{pk},3\,\rm{GHz}}=434\pm 21\,\rm{mJy}$.
The radio spectral index of our first late time ($\sim 80$ days post detection) ATCA measurement is $\alpha=0.58\pm 0.15$ where we use the convention $F_{\nu}\propto\nu^{-\alpha}$. Assuming this is associated with optically thin synchrotron emission where $\alpha=(p-1)/2$ this implies $p=2.16\pm 0.3$ where $p$ is the power-law index of the electron energy distribution ($N(E)\propto E^{-p}$). This indicates that the synchrotron cooling frequency, $\nu_{c}$, is above $9\,\rm{GHz}$ at this epoch, as otherwise a spectral slope of $-p/2$ would be expected and imply an extremely hard electron energy distribution. Our second ATCA epoch suggests a marginal steepening of the spectral index to $0.86\pm 0.30$, which would imply $p=2.72\pm 0.6$. In addition to the spectral index we can make crude estimates for the rise and decay indices of the $3\,\rm{GHz}$ light curves to be $\beta\sim-0.4$ and $\beta\sim 0.5$, respectively, for $F_{3\,\rm{GHz}}\propto t^{-\beta}$, although these calculations are obviously sensitive to the exact peak location and assume a constant evolution rate. The late time $3\,\rm{GHz}$ VLASS measurement indicates that no significant re-brightening occurred at late time.
@@ 4.1 Basic considerations
We begin with a general discussion of the radio properties of EP240414a in the context of other classes of extragalactic radio transients, particularly (relativistic) supernovae, fast blue optical transients, tidal disruption events (both thermal and relativistic), and gamma-ray bursts both long and short (these comparisons can be appreciated graphically by referring to Figure 2). The observed peak 3 GHz specific luminosity of EP240414a is $\gtrsim 3\times 10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and occurs at $T\sim T_{0}+20\,\rm{d}$ in the rest frame of the explosion. The timing of this peak is significantly earlier than for the radio counterparts to the luminous fast X-ray transients (LFBOTs), which typically peak closer to 100 days post explosion in the GHz range and at luminosities closer to $10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ (e.g. Margutti et al., 2019; Ho et al., 2019, 2020; Bright et al., 2022; Coppejans et al., 2020). An exception to this is the FBOT ZTF19abvkwla, which peaked close to $10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and at an uncertain peak time before 100 days post discovery Ho et al. (2020). This can be seen as highest luminosity green line in Figure 2.
@@
The peak luminosity of EP240414a also exceeds both regular and relativistic core-collapse supernovae by at least an order of magnitude. In a survey of almost 300 supernovae (including a range of different classifications) Bietenholz et al. (2021) found a peak specific radio luminosity of $10^{25.5\pm 1.6}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ (including non-detections), with the brightest source in the entire sample having a peak specific luminosity of $\sim 10^{29}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$. This is also consistent with the peak luminosity of radio emission from superluminous supernovae (Margutti et al., 2023), although radio detections of SLSNe are uncommon.
@@
Tidal disruption events (TDEs) are also know to produce radio emission, either from a relativistic jet or from non-relativistic outflow (e.g. van Velzen et al., 2016; Bright et al., 2018; Eftekhari et al., 2018; Alexander et al., 2020; Rhodes et al., 2023). These are termed ‘relativistic’ and ‘thermal’ TDEs, respectively. Relativistic TDEs are some of the most luminous radio transients known, with peak specific luminosities exceeding $\sim 10^{32}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and peaking at 100s of days post explosion (although the sample of such events is still relatively small). Thermal TDEs peak on similarly long timescales, but at much lower specific luminosities of $\sim 10^{29}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$. These peak luminosities and timescales are incompatible with those seen from EP240414a. Any potential TDE association for EP240414a is further disfavoured due to the significant projected offset ($\sim 27\,\rm{kpc}$) from the centre of the putative host galaxy (Jonker et al., 2024; Levan et al., 2024a, Srivastav et al. 2024; van Dalen et al. 2024) and see Figure 1.
@@
Finally, we consider EP240414a in the context of gamma-ray bursts (GRBs). Long GRB radio afterglows are caused by two major shocks, the forward and reverse, with radiation from each shock peaking between days and 100s of days post burst, depending on the observing frequency (see e.g. Laskar et al., 2013; van der Horst et al., 2014; Bright et al., 2023, for examples from well sampled long GRBs with two clear shock components). The peak radio luminosities of long GRBs span a wide range but are typically between $10^{30}$ and $\sim 10^{32}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ at $\sim 10$ days post discovery (see e.g. figure 6 of Chandra & Frail, 2012). The peak luminosity and timescale of EP240414a are both consistent with originating from a long GRB afterglow. The most significant evidence against a GRB progenitor is the location of EP240414a within its host galaxy. Numerous studies (Bloom et al., 2002; Blanchard et al., 2016; Lyman et al., 2017) have shown that long GRBs typically occur close to the centre of their host galaxy, with 80% of the sample presented in Lyman et al. (2017) located within $\sim 3\,\rm{kpc}$. In fact, the galactic offset of EP240414a ($\sim 27\,\rm{kpc}$) is much more consistent with those seen for short Gamma-ray bursts, with Fong et al. (2022) finding that a significant fraction of their sample fell outside of a $10\,\rm{kpc}$ radius and the most distant above $50\,\rm{kpc}$ in separation. However, EP240414a has been associated with a type Ic broad line supernovae (Levan et al. 2024a; van Dalen et al. 2024 in prep.), a supernova class solidly associated with long gamma-ray bursts (see Hjorth & Bloom, 2012, for a review). While luminosity rise-time arguments are not conclusive in determining the progenitor of a source, based on the arguments above we suggest that EP240414a is likely a long gamma-ray burst that occurred in an unusual location in its host galaxy. The lack of reported gamma-ray emission suggests that this GRB could have had a low isotropic gamma-ray luminosity, or the radio emission could be seen from off-axis. For example, the lack of detection by Konus-Wind suggests that the isotropic equivalent gamma-ray energy was below $\sim 10^{51}\,\rm{erg}$, based on the sample presented in Tsvetkova et al. (2017) (their figure 8). This energy is at the low end of the distribution for long GRBs, but not unusually so (see figure 1 in Perley et al., 2014).

### v2

@@ Abstract
Despite being operational for only a short time, the Einstein Probe mission, with its large field of view and rapid localisation capabilities, has already significantly advanced the study of rapid variability in the soft X-ray sky. We report the discovery of luminous and variable radio emission from the Einstein Probe fast X-ray transient EP240414a, the second such source with a radio counterpart. The radio emission at $3\,\rm{GHz}$ peaks at $\sim 30$ days post explosion and with a spectral luminosity $\sim 2\times 10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$, similar to what is seen from long gamma-ray bursts, and distinct from other extra-galactic transients including supernovae and tidal disruption events, although we cannot completely rule out emission from engine driven stellar explosions e.g. the fast blue optical transients. An equipartition analysis of our radio data reveals that an outflow with at least a moderate bulk Lorentz factor ($\Gamma\gtrsim 1.6$) with a minimum energy of $\sim 10^{48}\,\rm{erg}$ is required to explain our observations. The apparent lack of reported gamma-ray counterpart to EP240414a could suggest that an off-axis or choked jet could be responsible for the radio emission, although a low luminosity gamma-ray burst may have gone undetected. Our observations are consistent with the hypothesis that a significant fraction of extragalactic fast X-ray transients are associated with the deaths of massive stars.
@@ 1 Introduction
The observational paradigm for FXTs has recently undergone a drastic shift with the launch of the Einstein Probe X-ray telescope (EP; Yuan et al., 2022). The two instruments on board the EP, the Wide Field X-ray Telescope (WXT) and the Follow-up X-ray Telescope (somewhat unfortunately abbreviated as FXT), allow for both wide sky area monitoring as well as follow-up and localisation regions that range between arcminutes and tens of arcseconds, with FXT candidates reported via public alert streams. These capabilities have allowed for rapid follow-up at optical, radio, and X-ray wavelengths and the detection of multi-wavelength counterparts to EPW20240219aa, EP240305a, EP240309a, and EP240315a. The sources EP240305a and EP240309a are likely Galactic, and have been associated with variability from a Gaia star (Liu et al., 2024a) and a cataclysmic variable (Buckley et al., 2024), respectively. EPW20240219aa was associated with a sub-threshold event by the Fermi Gamma-ray Burst Monitor and is therefore thought to have been caused by a gamma-ray burst (Fletcher et al., 2024). EP240315a was associated with a high redshift ($z\sim 4.9$) galaxy and was followed up across the electromagnetic spectrum, being the first EP FXT with a radio and optical counterpart (Carotenuto et al., 2024; Gillanders et al., 2024). Analysis of the radio, optical, and X-ray data lead to the conclusion that the event harboured a relativistic jet, and was likely from a long gamma-ray burst (Gillanders et al., 2024; Ricci et al., 2024; Levan et al., 2024b; Liu et al., 2024b). With these events, and the increasing number of new events being discovered discovered and quickly reported by the EP, it is clear that significant progress can now be made in understanding the progenitors to FXTs and their multiwavelength properties.
@@
In this paper we present a radio observing campaign on EP240414a, the second EP extragalactic FXT with a multi-wavelength counterpart. EP240414a was discovered (Lian et al., 2024) by the WXT at UTC 09:50:12 2024-04-14 (MJD 60414.4099; which we define to be $T_{0}$) with a peak flux of $\sim 3\times 10^{-9}\,\rm{erg}\,\rm{s}^{-1}\,\rm{cm}^{-2}$ in the $0.5$-$4\,\rm{keV}$ energy band and was seen to fade by 4 dex over the following 2 hours with the EP Follow-up X-ray Telescope (Guan et al., 2024). An optical counterpart (AT2024gsa) was discovered within three hours of the FXT detection (Aryan et al., 2024) and a redshift of the nearest galaxy was reported to be $z=0.41$ (Jonker et al., 2024). The rapid and unusual optical lightcurve of the source is reported in Srivastav et al. (2024) with the redshift confirmed to be $z=0.4018\pm 0.0010$ (see also van Dalen et al., 2024). This implies a peak X-ray luminosity of $\sim 10^{48}\,\rm{erg}\,\rm{s}^{-1}$, similar to those seen from GRB afterglows at early times (see e.g. Gehrels et al., 2009), at a projected offset of $\sim 27\,\rm{kpc}$ from the host (unusually large for a long gamma-ray burst Bloom et al., 2002; Blanchard et al., 2016; Lyman et al., 2017). We discovered the radio counterpart to EP240414a with the MeerKAT radio interferometer, six days after it was reported as an FXT by the EP, as a $\sim 200\,\mu\rm{Jy}$ radio source coincident with the EP error circles (Guan et al., 2024; Lian et al., 2024) and coincident with AT2024gsa (Bright et al., 2024; Aryan et al., 2024; Srivastav et al., 2024; van Dalen et al., 2024).
@@ 2.1 MeerKAT
Figure 1: Optical and radio detections of EP240414a/AT2024gsa. (top) A Pan-STARRS i-band image of the field of EP240414a. The source position is marked by a pair of white lines. The diffuse source to the left is the putative host galaxy SDSS J124601.99-094309.3 (Aryan et al., 2024; Jonker et al., 2024) at a redshift of $z=0.4018\pm 0.0010$ (Srivastav et al. (2024), see also van Dalen et al. (2024)). (bottom) A subsection of our MeerKAT discovery image of EP240414a with the source position marked by a pair of black lines. The source has a flux density of $227\pm 13\,\mu\rm{Jy}$. We detect clear emission from the possible host galaxy of EP240414a. The MeerKAT restoring beam is shown in the bottom left and is $3.2^{\prime\prime}\times 3.2^{\prime\prime}$ at a position angle of 0 degrees. A scale bar in the bottom right shows $10\,\rm{kpc}$ at $z=0.4$.
@@ 2.2 Australia Telescope Compact Array
We obtained radio observations of EP240414a with the Australia Telescope Compact Array (ATCA) under program CX576 (PI Carotenuto). We first observed EP240414a on 2024 July 2 between 06:10 UT and 10:11 UT. ATCA was in its extended 6D configuration. A second observation, under the same program, was performed on August 23rd 2024 between 04:40 and 08:30 UT, with the telescope in the more compact 1.5A configuration. For both epochs, data were recorded simultaneously at central frequencies of 5.5 GHz and 9.0 GHz, with 2 GHz of bandwidth at each frequency. We used PKS B1934-638 for bandpass and flux density calibration, and J1239–1023 for the complex gain calibration. Data were flagged, calibrated, and imaged using standard procedures within CASA. When imaging, we used a Briggs robust parameter of 0 to balance sensitivity and resolution. Our ATCA observations are summarised in Table 1.
@@ 2.3 Allen Telescope Array
As part of a larger survey of EP transients we began observing the field of EP240414a on 2024 May 5 with the Allen Telescope Array (ATA; see Farah et al. in prep.; Pollak et al. in prep.). While multiple observing bands were used (see e.g. Bright et al., 2022, for details) we report only our most constraining limit at a central frequency of $8\,\rm{GHz}$. Observations were reduced with a custom pipeline utilising CASA for calibration and wsclean for imaging. 3C286 was used to set the absolute flux scale and bandpass response of the instrument, while 1246-075 was used to correct for the time-dependent complex gains. Our ATA observations are summarised in Table 1.
@@ 3 Results
Our radio observations of EP240414a are fairly sparse, with three observations at $3\,\rm{GHz}$ with MeerKAT, and two ATCA observations both at $5.5$ and $9\,\rm{GHz}$ (see Table 1 and Figure 2). Our $3\,\rm{GHz}$ light curve shows a clear peak at around $\sim 30$ days post explosion in the observer reference frame (or $\sim 20$ days in the source rest frame) at $434\pm 21\,\mu\rm{Jy}$. While our sampling rate is not high enough to confidently confirm this as the exact peak time at $3\,\rm{GHz}$, we can make the approximation that a spectral peak moved through $3\,\rm{GHz}$ at $30$ days post explosion and therefore take $T_{\rm{pk},3\,\rm{GHz}}=30$ days and $F_{\rm{pk},3\,\rm{GHz}}=434\pm 21\,\mu\rm{Jy}$.
The radio spectral index of our first late time ($\sim 80$ days post detection) ATCA measurement is $\alpha=0.58\pm 0.15$ where we use the convention $F_{\nu}\propto\nu^{-\alpha}$ and fit a simple power law between the two ATCA frequencies. Assuming this is associated with optically thin synchrotron emission where $\alpha=(p-1)/2$ this implies $p=2.16\pm 0.3$ where $p$ is the power-law index of the electron energy distribution ($N(E)\propto E^{-p}$). This indicates that the synchrotron cooling frequency, $\nu_{c}$, is above $9\,\rm{GHz}$ at this epoch, as otherwise a spectral slope of p/2-p/2 would be expected and imply an extremely hard electron energy distribution. Our second ATCA epoch suggests a marginal steepening of the spectral index to $0.86\pm 0.30$, which would imply $p=2.72\pm 0.6$, although the values are statistically consistent. In addition to the spectral index we can make crude estimates for the rise and decay indices of the $3\,\rm{GHz}$ light curves to be $\beta\sim-0.4$ and $\beta\sim 0.5$, respectively, for $F_{3\,\rm{GHz}}\propto t^{-\beta}$, although these calculations are obviously sensitive to the exact peak location and assume a constant evolution rate. The late time $3\,\rm{GHz}$ VLASS measurement indicates that no significant re-brightening occurred at late time.
@@ 4.1 Basic considerations
We begin with a general discussion of the radio properties of EP240414a in the context of other classes of extragalactic radio transients, particularly (relativistic) supernovae, fast blue optical transients, tidal disruption events (both thermal and relativistic), and gamma-ray bursts both long and short (these comparisons can be appreciated graphically by referring to Figure 2). The observed peak 3 GHz specific luminosity of EP240414a is $\gtrsim 3\times 10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and occurs at $T\sim T_{0}+20\,\rm{d}$ in the rest frame of the explosion. The timing of this peak is significantly earlier than for the radio counterparts to the luminous fast X-ray transients (LFBOTs), which typically peak closer to 100 days post explosion in the GHz range and at luminosities closer to $10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ (e.g. Margutti et al., 2019; Ho et al., 2019, 2020; Bright et al., 2022; Coppejans et al., 2020). An exception to this is the FBOT ZTF19abvkwla, which peaked close to $10^{30}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and at an uncertain peak time before 100 days post discovery (Ho et al., 2020). This can be seen as highest luminosity green line in Figure 2 and makes the association between FXTs and FBOTs hard to rule out.
@@
The peak luminosity of EP240414a also exceeds both regular and relativistic core-collapse supernovae by at least an order of magnitude. In a survey of almost 300 supernovae (including a range of different classifications) Bietenholz et al. (2021) found a peak specific radio luminosity of $10^{25.5\pm 1.6}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ (including non-detections), with the brightest source in the entire sample having a peak specific luminosity of $\sim 10^{29}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$. EP240414a is also more luminous at radio frequencies than all radio-detected superluminous supernovae (SLSN; Margutti et al., 2023; Eftekhari et al., 2021), although radio detections of SLSNe are uncommon, especially at early times.
@@
Tidal disruption events (TDEs) are also know to produce radio emission, either from a relativistic jet or from non-relativistic outflow (e.g. van Velzen et al., 2016; Bright et al., 2018; Eftekhari et al., 2018; Alexander et al., 2020; Rhodes et al., 2023). These are termed ‘relativistic’ and ‘thermal’ TDEs, respectively. Relativistic TDEs are some of the most luminous radio transients known, with peak specific luminosities exceeding $\sim 10^{32}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ and peaking at 100s of days post explosion (although the sample of such events is still relatively small). Thermal TDEs peak on similarly long timescales, but at much lower specific luminosities of $\sim 10^{29}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$. These peak luminosities and timescales are incompatible with those seen from EP240414a. Any potential TDE association for EP240414a is further disfavoured due to the significant projected offset ($\sim 27\,\rm{kpc}$) from the centre of the putative host galaxy (Jonker et al., 2024; Levan et al., 2024a; Srivastav et al., 2024; van Dalen et al., 2024) and see Figure 1.
@@
Finally, we consider EP240414a in the context of gamma-ray bursts (GRBs). Long GRB radio afterglows are caused by two major shocks, the forward and reverse, with radiation from each shock peaking between days and 100s of days post burst, depending on the observing frequency (see e.g. Laskar et al., 2013; van der Horst et al., 2014; Bright et al., 2023, for examples from well sampled long GRBs with two clear shock components). The peak radio luminosities of long GRBs span a wide range but are typically between $10^{30}$ and $\sim 10^{32}\,\rm{erg}\,\rm{s}^{-1}\,\rm{Hz}^{-1}$ at $\sim 10$ days post discovery (see e.g. figure 6 of Chandra & Frail, 2012). The peak luminosity and timescale of EP240414a are both consistent with originating from a long GRB afterglow. The most significant evidence against a GRB progenitor is the location of EP240414a within its host galaxy. Numerous studies (Bloom et al., 2002; Blanchard et al., 2016; Lyman et al., 2017) have shown that long GRBs typically occur close to the centre of their host galaxy, with 80% of the sample presented in Lyman et al. (2017) located within $\sim 3\,\rm{kpc}$. In the sample presented in Blanchard et al. 2016 there were no offsets above $20\,\rm{kpc}$ in their sample of $\sim 100$ LGBRs. In fact, the galactic offset of EP240414a ($\sim 27\,\rm{kpc}$) is much more consistent with those seen for short Gamma-ray bursts, with Fong et al. (2022) finding that a significant fraction of their sample fell outside of a $10\,\rm{kpc}$ radius and the most distant above $50\,\rm{kpc}$ in separation. However, EP240414a has been associated with a type Ic broad line supernovae (Levan et al. 2024a; van Dalen et al. 2024), a supernova class solidly associated with long gamma-ray bursts (see Hjorth & Bloom, 2012, for a review). The offset discrepancy was also commented on by van Dalen et al. 2024, where they note that luminous FBOTs can have host offsets comparable to EP240414a and that there is only weak star formation at the location of EP240414a in its host galaxy. While luminosity rise-time arguments are not conclusive in determining the progenitor of a source, based on the arguments above we suggest that EP240414a is likely a long gamma-ray burst that occurred in an unusual location in its host galaxy. The lack of reported gamma-ray emission suggests that this GRB could have had a low isotropic gamma-ray luminosity, or the radio emission could have been seen from off-axis. For example, the lack of detection by Konus-Wind suggests that the isotropic equivalent gamma-ray energy was below $\sim 10^{51}\,\rm{erg}$, based on the sample presented in Tsvetkova et al. (2017) (their figure 8). This energy is at the low end of the distribution for long GRBs, but not unusually so (see figure 1 in Perley et al., 2014). Following up sources detected with the Einstein Probe could help elucidate the distinction, or lack thereof (Nakar, 2015), between low luminosity LGRBs and regular LGRBs, potentially driven by differences in circumburst material or viewing angle (Liang et al., 2007; Salafia et al., 2016).